An 19.5 W device has 9.30 V across it. How much charge goes through the device in 4.15 h?
1
Expert's answer
2018-03-26T10:10:07-0400
The power
P=IV
So, the current through device
I=P/V=19.5/9.30=2.01 A
The amount of charge goes through the device
q=It=2.01×4.15×3600=3.00×〖10〗^4 C=30.0 kC